when 2 256 is divide by 17 the remainder would be?

 when 2256 is divide by 17 the remainder would be?

Grade:12

1 Answers

phani raaj
13 years ago

hey neha,

2^256 = 256^32

256^32 = (1+255)^32

apply binomial theorm,

(1+255)^32 = 1 + 255(k)

since 255 is divisible by 17

remainder will be 1